Show that if P and Q are vector Subspaces of a vector space V then P ∩ Q is also a vector subspace of V.


1
Expert's answer
2022-01-12T18:50:36-0500

Let "x,y\in P\cap Q" and "\alpha \in \mathbb{F}". We shall show that "P\cap Q\neq \phi" "\alpha(x+y) \in P\cap Q"

"\text{Since P and Q are subspaces of V then, P }≠\phi \text{ and } Q≠\phi\implies P\cap Q\neq \phi"

"\text{Since P and Q are subspaces of V }\implies \alpha(x+y) \in P \text{ and } \alpha(x+y) \in Q"

Thus, "\alpha(x+y) \in P\cap Q"

Hence, "P\cap Q" is a subspace of "V"